How many quarts of water [0% alcohol] need to be added to 50 quarts of 80% alcohol to obtain a 10% alcohol solution?
Found 2 solutions by stanbon, lwsshak3:
Answer by stanbon(75887)

How many quarts of water [0% alcohol] need to be added to 50 quarts of 80% alcohol to obtain a 10% alcohol solution?
----
Equation:
water + water = water
x + 0.20*50 = 0.90(x+50)
---
Multiply thru by 100 to get:
100x + 20*50 = 90x + 90*50
---
10x = 70*50
x = 350 quarts (amt. of water needed)
========================================
Cheers,
Stan H.

Answer by lwsshak3(11628)

How many quarts of water [0% alcohol] need to be added to 50 quarts of 80% alcohol to obtain a 10% alcohol solution?
**
let x=quarts of water to be added
80%*50+0%x=10%(50+x)
40+0=5+.1x
40-5=.1x
35=.1x
x=350 quarts
ans:
350 quarts of water need to be added to 50 quarts of 80% alcohol to obtain a 10% alcohol solution?